OpenStudy (anonymous):

find unit vectors parallel to tangent line to curve y=x^2 at point (2,4)

OpenStudy (anonymous):

Can you find f'(x) ?

OpenStudy (anonymous):

-> m = f'(2)

OpenStudy (anonymous):

@dizliz24 Hello...oo

OpenStudy (anonymous):

Sorry, I was working on the problem. I am completely lost on this one. I have no idea where to start

OpenStudy (anonymous):

I forget how to do derivatives

OpenStudy (anonymous):

f(x) = x^n --> f'(x) = nx ^(n-1) Now do you remember and start?

OpenStudy (anonymous):

Aha, yes. But how do I relate this to the information that I have available

OpenStudy (anonymous):

f(x) = x^n = x² Does it look related now?

OpenStudy (anonymous):

Yes, so I have 2x=4, correct? What should be my next step

OpenStudy (anonymous):

Do you understand what's 4?

OpenStudy (anonymous):

I am in a toss up. I am not sure if it is y or if it is 2x such that x=2 which is also 4

OpenStudy (anonymous):

Did you read my second post?

OpenStudy (anonymous):

yes. what is m

OpenStudy (anonymous):

y = mx + b What's m?

OpenStudy (anonymous):

I know that has something to do with slope

OpenStudy (anonymous):

Yes, slope! m is slope!

OpenStudy (anonymous):

Okay, m is the slope. So why does m=f'(2)

OpenStudy (anonymous):

That's how you find slope :)

OpenStudy (anonymous):

That's news to me. Now, what do I do with the slope

OpenStudy (anonymous):

So that's the basic you should memorize!

OpenStudy (anonymous):

Then what do you know about unit vector?

OpenStudy (anonymous):

Oh a unit vector is a vector whose length is 1

OpenStudy (anonymous):

How to find magnitude of vector?

OpenStudy (anonymous):

Thats one I am not sure of

OpenStudy (anonymous):

Do you know distance formula?

OpenStudy (anonymous):

d=sqrt((x2-x1)+(y2-y1)

OpenStudy (anonymous):

You miss the square sign (x2-x1)² + ( )² Great, distance formula to calculate magnitude of vector :)

OpenStudy (anonymous):

Yes, I did.

OpenStudy (anonymous):

How do I use this formula

OpenStudy (anonymous):

Since the unit vector parallel to m = 4 => it's coordinate < 1, 4> and it's magnitude √1 + 16 = √17 Thus the unit vector is < 1/√17, 4/√17 >

OpenStudy (anonymous):

Okay but why do we use 1, because it is a unit vector?

OpenStudy (anonymous):

When you divide for magnitude => Unit vector!

OpenStudy (anonymous):

what do you divide

OpenStudy (anonymous):

I don't understand what we are dividing and why we divide

OpenStudy (anonymous):

Then you need to look at your note, notebook about unit vector!

OpenStudy (anonymous):

I see what you mean. I feel stupid. thank you for your help:)